Anderson's supploers extend credit to the frim on terms of "net 30." Anderson normally pays its bills on the last day of the dredit period. How much additional financing could be generated if Anderson were to stretch its payables 10 days beyond due date?TABLE 160.1 Anderson Furniture Companys Financi urniture Companys Financial Data (in Thousands of Dollars) Industry Average
Answer #1

Answer:

Current credit period given by suppliers = 30 days

We observe from Anderson's balance sheet:

Accounts Payable = $3,000,000

Hence:

If Anderson were to stretch its payable 10 days beyond due date, new average payable period = 30 + 10 = 40

New average accounts payable = 3,000,000/ 30 * 40 = $4,000,000

Additional financing that could be generated = 4000000 - 3000000 = $1,000,000

Additional financing could be generated if Anderson were to stretch its payable by 10 days = $1,000,000